Given the equation (x - 3)2 + y2 = 121 of a circle, identify the center and the radius.
Answer by Edwin McCravy(20086)



(x - 3)² + y² = 121

Ww want to first write it in the form 

(x - h)² + (y - k)² = r², which has center (h,k) and radius r

(x - 3)² + y² = 121

The first term is already in the form (x - h)².  The second
term is y², so to get it in the form (y - k)², we write it 
as (y - 0)²

(x - 3)² + (y - 0)² = 121

To get the last term in the form r² we write 121 as 11²:

(x - 3)² + (y - 0)² = 11²

Now we compare it to 

(x - h)² + (y - k)² = r²

and get h = 3, k = 0, r = 11

So the center is (h,k) = (3,0) and the radius is r = 11

Edwin